OpenStudy (love_to_love_you):

PLEASE HELP WILL FAN AND MEDAL Write a polynomial function with the given zeros. Show all work. a. x= -1, 0, 2 b. x= -2, 1, 4

ganeshie8 (ganeshie8):

heard of factor theorem ?

ganeshie8 (ganeshie8):

If \(x=k\) is a zero of polynomial then \((x-k)\) is a factor

OpenStudy (love_to_love_you):

No

OpenStudy (love_to_love_you):

Never mind I just realized I asked this already haha

OpenStudy (love_to_love_you):

if x = -1 is a zero or root, then (x + 1) is a factor x = 2 is zero or root then (x -2) is a factor x = 0 is a zero then x is a factor P(x) =x(x-2)(x+1) F(x)= x^3-x^2-2x